Solve this ∫(xsinx/e x +1) dx Limit is from -1 to 1.

Solve this
∫(xsinx/ex+1) dx
Limit is from -1 to 1.

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int_{-1}^{1}\frac{xsinx}{e^{x}+1}dx
I = \int_{0}^{1}(\frac{xsinx}{e^{x}+1}+\frac{xsinx}{e^{-x}+1})dx
I = \int_{0}^{1}xsinx(\frac{1}{e^{x}+1}+\frac{e^x}{e^{x}+1})dx
I = \int_{0}^{1}xsinxdx
I = (-xcosx+sinx)_{0}^{1}
I = (-cos1+sin1)
I = sin(1) - cos(1)